OpenStudy (rowanmartinfan):

mathway ppl ok so that gave me x^5 & y^2 its not an answer

OpenStudy (kainui):

What is the question?

OpenStudy (rowanmartinfan):

(x^2y)(x^4*y^3) --------------- xy^2

OpenStudy (rowanmartinfan):

the ^ are to show the # is a power

OpenStudy (kainui):

\[\frac{ (x^2y)(x^4y^3) }{ xy^2 }\]

OpenStudy (kainui):

is it like that?

OpenStudy (rowanmartinfan):

how u get the power up there

OpenStudy (kainui):

\[\frac{ (x^2y)(x^4y^3) }{ xy^2 } = x ^{4+2-1}y ^{3+1-2} = x^5y^2\]

OpenStudy (rowanmartinfan):

that is how it is ther

OpenStudy (rowanmartinfan):

thanks

OpenStudy (kainui):

You're welcome

OpenStudy (rowanmartinfan):

wanna help me with another one Kainui

OpenStudy (rowanmartinfan):

\[\frac{ x^6y^4 }{ x*y^2 }\]

OpenStudy (rowanmartinfan):

@Kainui \[\frac{ x^6y^4 }{ xy^2 }\]

OpenStudy (kainui):

\[x ^{6-1}y ^{4-2}?\]

OpenStudy (rowanmartinfan):

5 & 2

OpenStudy (kainui):

\[x^n = \frac{ 1 }{ x ^{-n} }\]

OpenStudy (kainui):

yes

OpenStudy (rowanmartinfan):

ok thanks

OpenStudy (rowanmartinfan):

@Kainui yay :D we got them right

OpenStudy (kainui):

:)

OpenStudy (kainui):

Glad I could help

OpenStudy (rowanmartinfan):

@Kainui thanks i might need more help tho again

OpenStudy (rowanmartinfan):

NOOOOO RAISING A PWR TO PWR

OpenStudy (kainui):

Just multiply

OpenStudy (rowanmartinfan):

oh so both pwrs together

OpenStudy (kainui):

can you start a new question, this one is lagging me since it got too long! Thanks!

OpenStudy (rowanmartinfan):

ooh ok will do